OpenStudy (anonymous):

what is tan of 105 degrees?

OpenStudy (anonymous):

Put in calculator tan(105)= -3.73

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

my guess is you are supposed to use the addition angle formula writing \(105=60+45\)

OpenStudy (misty1212):

\[\tan(60+45)=\frac{\tan(60)+\tan(45)}{1-\tan(60)\tan(45)}\] if i remember that formula correctly

OpenStudy (anonymous):

exactly -3.73205080757

OpenStudy (misty1212):

lol that is not "exact" dear, that is a decimal approximation a good one, but not "exact"

OpenStudy (anonymous):

(tan60+tan45)/(1-tan60 tan45) (radical3+1)/(1-radical3*1) (radical3+1)/(radical3-1) (radical3+1)^2/ (radical3-1) (radical3+1) (4+2radical3)(3-1) (2+radical3) = -3.732050807568877293527446341507

OpenStudy (anonymous):

100%